to earn an A in a course, a student must get at least a 90 average on four tests and a final exam, with the final exam weighted twice that of any one test. If the four test scores are: 93, 69, 89, and 97, what is the lowest score the student can earn on the final exam and still get an A
Answer by rfer(16322)

(93+69+89+97+x)/5=90
(348+x)/5=90
348+x=90*5
348+x=450
x=102